OpenStudy (anonymous):

Find a row echelon form fo the matrix 1 3 -1 2 1 4 -3 0 1 Please include the steps! thank you so much :)

OpenStudy (anonymous):

No, you do it.

OpenStudy (anonymous):

:(

hartnn (hartnn):

do you know what a 'row echelon form' is ?

OpenStudy (anonymous):

yes! like 1 01 001 format?

hartnn (hartnn):

thats correct, have you tried this ? if yes, which step are you stuck ?

OpenStudy (anonymous):

well... im not sure where to begin...it doesnt end right..

hartnn (hartnn):

ok, so firstly, you know that you want to make that '2' as 0, right ? for that you can do, R2 = R2-2R1 what u get ?

OpenStudy (anonymous):

wait.. do u mean -2R1 - R2?

hartnn (hartnn):

-2(1) -2 = -2-2 =-4 not that won't give you 0, right ?

OpenStudy (anonymous):

oops i mean +

hartnn (hartnn):

then that what i said...

OpenStudy (anonymous):

so 0 -5 6?

hartnn (hartnn):

yes, thats correct.

hartnn (hartnn):

now how about making that '-3' as 0 ? do, R3=3R1+R3

OpenStudy (anonymous):

0 9 -2

hartnn (hartnn):

so now, you only require o get rid of that 9 , right ?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

idk how to do that because if i try to get rid of it the 0 would change?

OpenStudy (anonymous):

OHHH MULTIPLY ROW 2 BY 9/5

OpenStudy (anonymous):

5/9 *

OpenStudy (anonymous):

nvm

hartnn (hartnn):

MULTIPLY ROW 2 BY 9/5<---correct then you can add row 2 and row 3

OpenStudy (anonymous):

i got 0 0 44/5

hartnn (hartnn):

-2+6*9/5 = -2+54/5 = 44/5 yes, thats correct.

OpenStudy (anonymous):

now the 2nd number in row 2 has to be 1 right?

hartnn (hartnn):

is it required ? http://en.wikipedia.org/wiki/Row_echelon_form not actually. you're done.

hartnn (hartnn):

if you were asked 'reduced' row ...then you need to make that =1

OpenStudy (anonymous):

OHHHHH OKAY I GET IT!

OpenStudy (anonymous):

THANK YOU SOOO MUCH :)

hartnn (hartnn):

welcome ^_^ i see you are new here, so \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

:)))) thank you!!!!